) Let T be a normal operator on a finite dimensional Hilbert space H

with spectrumπœ†1, πœ†2, … … , πœ†π‘š , . Then prove that

i) T is self - adjoint⟺ each πœ†π‘–,,is real

ii) T is positive ⟺ each πœ†π‘– β‰₯ 0

iii) T is unitary ⟺ πœ†π‘– =1 for each .


1
Expert's answer
2021-12-01T12:28:19-0500

A complex number λ ∈ C is called an eigenvalue of T ∈ B(H) if there exists a vector 0 "\neq" x ∈ H such that Tx = λx


i)

Let Ξ» be an eigenvalue of T and x be the corresponding eigenvector. Then Tx = Ξ»x.

"\lambda||x||^ 2 = λ \langle x,x\rangle = \langle λ x,x\rangle = \langle T x,x\rangle = \langle x,Tx\rangle= \langle x, λ x\rangle =\overline{\lambda}||x||^ 2"

Since x "\neq" 0, we have Ξ» = "\overline{ λ}"


ii)

T is positive if "Tx\ge 0"

then Tx = Ξ»x "\ge 0 \implies \lambda \ge 0"


iii)

"Tx=T^{-1}x"

"\lambda x=\lambda^{-1} x \implies \lambda=1"